# Nemotron-CC: Transforming Common Crawl into a Refined Long-Horizon Pretraining Dataset

### v1

@@ 1 Introduction
    We propose a method for transforming English Common Crawl into a 6.3T token long-horizon pretraining dataset, consisting of 4.4T globally deduplicated original tokens and 1.9T synthetically generated tokens. We release the dataset [^2] https://data.commoncrawl.org/contrib/Nemotron/Nemotron-CC/index.html and plan to release an implementation as part of the open-source NeMo Curator library. [^3] https://github.com/NVIDIA/NeMo-Curator
@@ 2 Methods
In this section we explain our efforts to build the best English Common Crawl pretraining dataset for LLMs. Our efforts can be splitted into three folds. First, we talk about our efforts in boosting token yield by utilizing text extractor and heuristic filters more properly in Section 2.1. Second, we introduce the model-based quality labeling pipeline methods in Section 2.2. Third, we introduce our synthetic data generation method to further improve the data quality in Section 2.3.
@@ 2.1 HTML-to-text Extractor & Filter
We test two HTML-to-text extractors, Justext (Pomikálek, 2011) and Trafilatura (Barbaresi, 2021). Qualitatively, we view both extractors at the same level of quality. While quantitatively, we calculate token yields of both extractors on 13 selected snapshots of Common Crawl (see Appendix D). The statistics were reported in Table 1. We see that Justext can yield more tokens, notably more high-quality tokens (+28.6%) by the standard of Fineweb-Edu classifier (score 3, 4, and 5). We highlight that, boosting unique token amount is of great importance when building long-horizon pretraining dataset, e.g., 15T tokens for Llama3.1. We keep only English text, as determined by pycld2444https://pypi.org/project/pycld2/ and the FastText lid176 language classifier [^5] https://fasttext.cc/docs/en/language-identification.html with threshold 0.3 (Joulin et al., 2016a, b). After extraction, we apply global fuzzy deduplication as well as exact substring deduplication (Lee et al., 2022), using the NeMo Curator library [^6] https://github.com/NVIDIA/NeMo-Curator and the deduplicate-text-datasets library, [^7] https://github.com/google-research/deduplicate-text-datasets respectively.
@@ 2.2 Model-based Quality Labeling
Recent work (Li et al., 2024; Penedo et al., 2024) use model-based classifiers to extract high-quality pretraining documents from English Common Crawl. However, both of the two quality classifiers have a limited recall (less than 10%) of high-quality tokens, and this will become a bottleneck to train an LLM over a long horizon. Also, the quality labels assigned by the quality classifier are not necessarily aligned with LLM’s downstream task performance. Therefore, we propose our ensemble-based quality labeling pipeline method. Specifically, we first build three quality classifiers, each of which has different high-quality preferences. Then, we ensemble the three classifiers to score all the documents, and split the crawl corpus into different quality buckets based on the quality score. Finally, we regroup the fine-grained document buckets into 5 different quality levels based on their corresponding performance on downstream task.
@@
In order to assign a quality label that is more aligned with their real performance on downstream tasks, we further group the fine-grained quality score predicted by three classifiers into 5 downstream quality categories. We used annealing to access each data bucket’s downstream task’s quality. Specifically, we measure the quality of each bucket by continuous pretraining with 50B tokens on a 70% trained 8B models. We assign 66% of weight to the default data mix and 34% to the dataset that we are evaluating. By comparing the average performance of each bucket over 9 tasks, we group the 20 buckets into 5 big categories, with the final distribution shown in Table 2.
@@
High 19 12.63
Medium-High 18 11.52
Medium 12-17 46.24
Medium-Low 7-11 20.43
Low 0-6 9.18
@@ 2.3 Synthetic Data Generation
For low-quality data, our goal is to improve the quality by reducing noise and errors while preserving useful information, thereby decreasing training compute expenses. As shown by Maini et al. (2024), rephrasing web data using a medium-sized language model yields an enhanced parallel corpus of synthetic data, thereby reducing model perplexity and boosting its accuracy on downstream tasks. Unlike existing methods that create new content such as textbooks and short stories Wang et al. (2023); Eldan and Li (2023); Gunasekar et al. (2023), our rephrasing-based approach does not utilize the language model as a knowledge bank but focuses on transforming provided texts into another style, allowing it to operate with a lighter-weight model. We adopt the Wikipedia style prompt from (Maini et al., 2024) to rewrite low-quality documents (Prompt 5 in Appendix F), which effectively reduces errors and redundancies and improves formatting.
For high-quality data, we aim to obtain more unique tokens and condense essential knowledge. According to (Muennighoff et al., 2024), adding repeated tokens yields a diminishing return, especially after 4 epochs. For high-quality documents, we generate synthetic data using four additional prompts: (1) Diverse Question-Answer (QA) pairs: ask questions in various forms (e.g., yes/no question, open-ended question, multi-choice question) about factual information in the text and provide the correct answers; (2) Distill: rewrite the text into a concise and clear passage; (3) Extract knowledge: rewrite knowledge from the text and disregard uninformative content; (4) Knowledge list: extract key information from the text as an organized list. We require the model to provide clear and concise responses while preserving factual information and concrete details such as numbers. The full prompts are shown in Appendix F.
As we increase the length of provided text, the model shows a tendency to produce over-simplified outputs with reduced detail. Therefore, we chunk each document into segments, each of which contains one or more complete lines and is shorter than a specific token limit. [^10] The token limit is set to 512 for Wikiepdia, 2,000 for Distill, 1,400 for Extract Knowledge and 1,000 for Diverse QA Pairs and Knowledge List, including tokens from the prompt and chat format. Over-length lines exceeding the token limit are discarded.
@@
Using the instruct version of Mistral NeMo 12B [^11] https://mistral.ai/news/mistral-nemo with FP8 inference, we synthesize over 1.8T tokens as Table 3 shows, including 336.3B tokens from low-quality documents and 1.5T tokens from high-quality documents. We do not use medium-quality documents for synthetic data generation. We employ TensorRT-LLM [^12] https://github.com/NVIDIA/TensorRT-LLM and NeMo-Skills [^13] https://github.com/NVIDIA/NeMo-Skills to enable large-scale data synthesis.
@@ 3.1 Experiment Setup
We use the open source Megatron-LM library [^14] https://github.com/NVIDIA/Megatron-LM (Shoeybi et al., 2019) to train standard 8B parameter transformer LLMs. The hyperparameter details are shown in Appendix C.
@@
Unless otherwise noted, we train for 1T tokens on a blend of $73\%$ English Common Crawl data and $27\%$ a fixed mix of specialized math, code, papers, books, patents, and Wikipedia datasets (Adler et al., 2024). When comparing datasets, we vary only the $73\%$ English Common Crawl portion.
@@ 3.2 Main Results
Our dataset contributed 7.2T of the tokens used to train an 8B model for 15T tokens. As shown in Table 6, our model achieves a higher average accuracy than Llama 3.1 8B, which was also trained for 15T tokens, including an MMLU score of 70.3 vs. Llama’s 65.3. This shows that our dataset is indeed suitable for state-of-the-art training over long token horizons.
@@ 3.3 Ablation Study
Justext HQ unfiltered 57.5 60.6
Table 7: Ablation studies on extractor and filter. HQ means high-quality data judged by FineWeb-Edu classifier (score 3,4,5). See Appendix E for more details.
@@
We did a detailed comparison of two types of classifiers that we employ in our method: the FineWeb-Edu classifier which score document quality based on their educational-level, and the DCLM-based classifier which value the informativeness of the document. We compare the high-quality documents predicted by the two classifiers on a randomly selected Common Crawl Snapshot (CC-MAIN-2021-21). Table 8 shows the document statistics comparison. We can see that only 10% of the documents are predicted as high quality by both classifiers, while 35.4% documents are predicted as high quality by FineWeb-Edu classifier only, and 54.4% of documents are predicted as high-quality by DCLM classifier. Therefore, ensembling different classifiers can increase the recall of high-quality documents from Common Crawl. [^16] Detailed URL domain comparison can be found in Appendix A
We further compare each of the classifiers with the ensembled method [^17] Note that we did not employ FineWeb-Edu classifier in our ensemble for license issue, since it is trained with annotations from Llama3. by their downstream tasks’ performances. We pretrain 8B parameters LLMs with 1T tokens, using the high-quality documents labeled by different classifiers on randomly selected 13 Common Crawl snapshots (see Appendix D). Table 9 shows the detailed comparison on different evaluation tasks. We can see that the ensembled method greatly boost the high-quality tokens percentage from 9% to 25%, while still achieving the highest general language understanding performance on average on all the tasks. The ensembled method also outperforms the FineWeb-Edu classifier and the DCLM classifier, in terms of the high-quality token percentage, and is on-par or slightly better on the 9 evaluation tasks. This is very important since more unique high-quality tokens is the key in pretraining larger LLMs on longer tokens horizons.
@@
As Table 10 shows, this ablation study aim to answer two questions: (1) Does rephrasing low-quality improve accuracies on downstream tasks? (2) Can synthetic data help offset the decreasing value of duplicated data reported in (Muennighoff et al., 2024)? To answer these questions, we train four 8B models with the same hyperparameters on different blends of 1T tokens: (1) LQ-Base: original Common Crawl data; (2) LQ-Synthetic: an augmented version of LQ-Base where the low-quality documents are rephrased; (3) HQ-Base: a blend containing eightfold high-quality documents and less low- and medium-quality documents; (4) HQ-Synthetic: a variant of HQ-Base where 4 repetitions of the high-quality documents are swapped out for synthetic datasets.
By comparing the results between LQ-Base and LQ-Synthetic, we can see that rephraing low-quality data leads to 1.50 absolute gains on average score. We also observe noticeable boosts from 1.80% to 4.75% on ARC-Easy, ARC-Challenge, OpenbookQA, CommonsenseQA; however, we also encounter slight accuracy drops on some tasks, which may indicate potential misinformation introduced by data synthesis. Current practices typically utilize data curation approaches to detect and eliminate noisy examples. Due to time and resource constraints, we leave the detailed exploration of this issue for future efforts.
@@ 4 Related Work
Synthetic datasets have been widely used in language model pre-training and post-training. In (Eldan and Li, 2023), the authors show that smaller or simpler models trained on a synthetic dataset of short stories are capable of generating fluent and consistent stories. Similarly, smaller models trained using high-quality synthetic textbook and exercise datasets can achieve impressive high accuracy on coding benchmarks (Gunasekar et al., 2023; Li et al., 2023). These approaches typically require a powerful language model, such as GPT-3.5 and GPT-4 in (Eldan and Li, 2023), to synthesize new contents. Instead, (Maini et al., 2024) shows that compact models such as Qwen-1.8B and Mistral-7B are adequate to rephrase web data. This approach generates diverse, high-quality synthetic data that effectively lowers model perplexity and boosts performance across benchmarks. We adopt this main idea, but explore more prompts and show how to specialize them for low and high quality data.
@@ 5 Conclusion
For producing long-horizon pretraining tokens for LLM from English Common Crawl data, we showed how to improve upon the state of the art and achieve better trade-offs between benchmark accuracy and data quantity, as measured by number of unique real tokens. Specifically, we showed the efficacy of ensembling model-based quality filters, rephrasing low and high quality documents, and reducing the reliance on non-learned heuristics.
@@ 6 Limitations
Some of the key limitations of our work are as follows. For the model-based filter ensembling and quality bucketing, we only had time and resources to try a single strategy. Though it is effective, it is possible this could be improved upon in future work, especially to improve the sensitivity at the higher-quality end of the spectrum. For the rephrased data, we did not verify the factual accuracy or fidelity to the original contents. More work is required to understand the risks of hallucinations or loss of content diversity in this setting and how to mitigate them. We also only looked at rephrasing low and high quality data. It could be interesting to explore how to best rephrase medium quality data as well. We did not do ablations on all parts of the pipeline. There is probably room for improvement with, for example, the language identification. Finally, we tried our methods only on English text. More work is needed to adapt our methods to other languages.

### v2

@@ 1 Introduction
    We propose a method for transforming English Common Crawl into a 6.3T token long-horizon pretraining dataset, consisting of 4.4T globally deduplicated original tokens and 1.9T synthetically generated tokens. We release the dataset [^2] https://data.commoncrawl.org/contrib/Nemotron/Nemotron-CC/index.html under the Common Crawl Terms of Use and a reference implementation as part of the Apache 2.0 open-source NeMo Curator library. [^3] https://github.com/NVIDIA/NeMo-Curator The quality classifier models have been released as well. [^4] https://huggingface.co/nvidia/nemocurator-fineweb-nemotron-4-edu-classifier and https://huggingface.co/nvidia/nemocurator-fineweb-mixtral-edu-classifier
@@ 2 Methods
In this section we explain our efforts to build the best English Common Crawl pretraining dataset for LLMs. Our efforts can be split into three folds. First, we talk about our efforts in boosting token yield by utilizing text extractor and heuristic filters more properly in Section 2.1. Second, we introduce the model-based quality labeling pipeline methods in Section 2.2. Third, we introduce our synthetic data generation method to further improve the data quality in Section 2.3. For a schematic overview of our final pipeline, please see Figure 3 in Appendix A.
@@ 2.1 HTML-to-text Extractor & Filter
We test two HTML-to-text extractors, Justext (Pomikálek, 2011) and Trafilatura (Barbaresi, 2021). Qualitatively, we view both extractors at the same level of quality. Quantitatively, we calculate token yields of both extractors on 13 selected snapshots of Common Crawl (see Appendix F). The statistics are reported in Table 1. We see that Justext can yield more tokens, notably more high-quality tokens (+28.6%) by the standard of Fineweb-Edu classifier (score 3, 4, and 5). We highlight that boosting unique token amount is of great importance when building long-horizon pretraining dataset, e.g., 15T tokens for Llama3.1. Even though there is a slight decline in the percentage of HQ tokens for Justext vs. Justext-filtered ($7.0\%$ vs. $7.5\%$), what we aim to maximize here is the absolute number of HQ tokens (127B vs. 104B). We will later sort the data into quality buckets, which enables exact control of the proportion of HQ vs. non-HQ data seen during training instead of reliance on the natural distribution for a particular extraction tool. After extraction, we apply filtering to keep only English text, as determined by pycld2555https://pypi.org/project/pycld2/ and the FastText lid176 language classifier [^6] https://fasttext.cc/docs/en/language-identification.html with threshold 0.3 (Joulin et al., 2016, 2017). We then apply global fuzzy deduplication as well as exact substring deduplication over eighths of snapshots (Lee et al., 2022), using the NeMo Curator library [^7] https://github.com/NVIDIA/NeMo-Curator and the deduplicate-text-datasets library, [^8] https://github.com/google-research/deduplicate-text-datasets respectively.
@@ 2.2 Model-based Quality Labeling
Recent work (Li et al., 2024; Penedo et al., 2024) use model-based classifiers to extract high-quality pretraining documents from English Common Crawl. However, both of the two quality classifiers have a limited recall (around 10%) of high-quality tokens (see Table 9), and this will become a bottleneck to train an LLM over a long horizon. Also, the quality labels assigned by the quality classifier are not necessarily aligned with LLM’s downstream task performance. Therefore, we propose our ensemble-based quality labeling pipeline method. Specifically, we first build three quality classifiers, each of which has different high-quality preferences. Then, we ensemble the three classifiers to score all the documents, and split the crawl corpus into different quality buckets based on the quality score. Finally, we regroup the fine-grained document buckets into 5 different quality levels based on their corresponding performance on downstream task.
@@
In order to assign a quality label that is more aligned with their real performance on downstream tasks, we further group the fine-grained quality score predicted by three classifiers into 5 downstream quality categories. We used annealing to assess each data bucket’s downstream task’s quality. Specifically, we measure the quality of each bucket by continuous pretraining with 50B tokens on a 70% trained 8B models. We assign 66% of weight to the default data mix and 34% to the dataset that we are evaluating. By comparing the average performance of each bucket over 9 tasks, we group the 20 buckets into 5 big categories, with the final distribution shown in Table 2. For more details, please see Appendix C.
@@
High 19 553 12.63
Medium-High 18 504 11.52
Medium 12-17 2,023 46.24
Medium-Low 7-11 894 20.43
Low 0-6 402 9.18
@@ 2.3 Synthetic Data Generation
For low-quality data, our goal is to improve the quality by reducing noise and errors while preserving useful information, thereby decreasing training compute expenses. As shown by Maini et al. (2024), rephrasing web data using a medium-sized language model yields an enhanced parallel corpus of synthetic data, thereby reducing model perplexity and boosting its accuracy on downstream tasks. Unlike existing methods that create new content such as textbooks and short stories Wang et al. (2023); Eldan and Li (2023); Gunasekar et al. (2023), our rephrasing-based approach does not utilize the language model as a knowledge bank but focuses on transforming provided texts into another style, allowing it to operate with a lighter-weight model. We adopt the Wikipedia style prompt from (Maini et al., 2024) to rewrite low-quality documents (Prompt 5 in Appendix H), which effectively reduces errors and redundancies and improves formatting.
For high-quality data, we aim to obtain more unique tokens and condense essential knowledge. According to (Muennighoff et al., 2024), adding repeated tokens yields a diminishing return, especially after 4 epochs. For high-quality documents, we generate synthetic data using four additional prompts: (1) Diverse Question-Answer (QA) pairs: ask questions in various forms (e.g., yes/no question, open-ended question, multi-choice question) about factual information in the text and provide the correct answers; (2) Distill: rewrite the text into a concise and clear passage; (3) Extract knowledge: rewrite knowledge from the text and disregard uninformative content; (4) Knowledge list: extract key information from the text as an organized list. We require the model to provide clear and concise responses while preserving factual information and concrete details such as numbers. The full prompts are shown in Appendix H.
As we increase the length of provided text, the model shows a tendency to produce over-simplified outputs with reduced detail. Therefore, we chunk each document into segments, each of which contains one or more complete lines and is shorter than a specific token limit. [^11] The token limit is set to 512 for Wikipedia, 2,000 for Distill, 1,400 for Extract Knowledge and 1,000 for Diverse QA Pairs and Knowledge List, including tokens from the prompt and chat format. Over-length lines exceeding the token limit are discarded.
@@
Using the instruct version of Mistral NeMo 12B [^12] https://mistral.ai/news/mistral-nemo with FP8 inference, a top-p value of $0.9$, and a sampling temperature of $0.5$, we synthesize over 1.8T tokens as Table 3 shows, including 336.3B tokens from low-quality documents and 1.5T tokens from high-quality documents. We do not use medium-quality documents for synthetic data generation due to time and resource constraints. We employ TensorRT-LLM [^13] https://github.com/NVIDIA/TensorRT-LLM and NeMo-Skills [^14] https://github.com/NVIDIA/NeMo-Skills to enable large-scale data synthesis.
@@ 3.1 Experiment Setup
We use the open source Megatron-LM library [^15] https://github.com/NVIDIA/Megatron-LM (Shoeybi et al., 2019) to train standard 8B parameter transformer LLMs. The hyperparameter details are shown in Appendix D.
@@
Unless otherwise noted, we train for 1T tokens on a blend of $73\%$ English Common Crawl data and $27\%$ a fixed mix of specialized code, papers, books, patents, and Wikipedia datasets (Adler et al., 2024). When comparing datasets, we vary only the $73\%$ English Common Crawl portion. See Table 12 in Appendix D.
@@ 3.2 Main Results
Our dataset contributed 7.2T of the tokens used to train an 8B model for 15T tokens. As shown in Table 6, our model achieves a higher average accuracy than Llama 3.1 8B, which was also trained for 15T tokens, including an MMLU score of 70.3 vs. Llama’s 65.3. This shows that our dataset is indeed suitable for state-of-the-art training over long token horizons. For more details on this experiment, please see Appendix E.
@@ 3.3 Ablation Study
Justext HQ-unfiltered 57.5 60.6
Table 7: Ablation studies on extractor and filter. HQ means high-quality data judged by FineWeb-Edu classifier (score 3,4,5). HQ-unfiltered means filtering is applied only to LQ data. See Appendix G for more details.
@@
We did a detailed comparison of two types of classifiers that we employ in our method: the FineWeb-Edu classifier which score document quality based on their educational-level, and the DCLM-based classifier which value the informativeness of the document. We compare the high-quality documents predicted by the two classifiers on a randomly selected Common Crawl Snapshot (CC-MAIN-2021-21). Table 8 shows the document statistics comparison. We can see that only 10% of the documents are predicted as high quality by both classifiers, while 35.4% documents are predicted as high quality by FineWeb-Edu classifier only, and 54.4% of documents are predicted as high-quality by DCLM classifier. Therefore, ensembling different classifiers can increase the recall of high-quality documents from Common Crawl. [^17] Detailed URL domain comparison can be found in Appendix B
We further compare each of the classifiers with the ensembled method [^18] Note that we did not employ FineWeb-Edu classifier in our ensemble for license issue, since it is trained with annotations from Llama3. by their downstream tasks’ performances. We pretrain 8B parameters LLMs with 1T tokens, using the high-quality documents labeled by different classifiers on randomly selected 13 Common Crawl snapshots (see Appendix F). Table 9 shows the detailed comparison on different evaluation tasks. We can see that the ensembled method greatly boost the high-quality tokens percentage from 9% to 25%, while still achieving the highest general language understanding performance on average on all the tasks. The ensembled method also outperforms the FineWeb-Edu classifier and the DCLM classifier, in terms of the high-quality token percentage, and is on-par or slightly better on the 9 evaluation tasks. This is very important since more unique high-quality tokens is the key in pretraining larger LLMs on longer tokens horizons.
@@
As Table 10 shows, this ablation study aim to answer two questions: (1) Does rephrasing low-quality improve accuracies on downstream tasks? (2) Can synthetic data help offset the decreasing value of duplicated data reported in (Muennighoff et al., 2024)? To answer these questions, we train four 8B models with the same hyperparameters on different blends of 1T tokens: (1) LQ-Base: original Common Crawl data including low-quality documents; (2) LQ-Synthetic: an augmented version of LQ-Base where the low-quality documents are rephrased; (3) HQ-Base: a blend containing eightfold high-quality documents and less low- and medium-quality documents; (4) HQ-Synthetic: a variant of HQ-Base where 4 repetitions of the high-quality documents are swapped out for synthetic datasets.
By comparing the results between LQ-Base and LQ-Synthetic, we can see that rephrasing low-quality data leads to 1.50 absolute gains on average score. We also observe noticeable boosts from 1.80% to 4.75% on ARC-Easy, ARC-Challenge, OpenbookQA, CommonsenseQA; however, we also encounter slight accuracy drops on some tasks, which may indicate potential misinformation introduced by data synthesis. Current practices typically utilize data curation approaches to detect and eliminate noisy examples. Due to time and resource constraints, we leave the detailed exploration of this issue for future efforts.
@@ 4 Related Work
Synthetic datasets have been widely used in language model pre-training and post-training. In (Cheng et al., 2024), instruction-response pairs are synthesized for pre-training. In (Eldan and Li, 2023), the authors show that smaller or simpler models trained on a synthetic dataset of short stories are capable of generating fluent and consistent stories. Similarly, smaller models trained using high-quality synthetic textbook and exercise datasets can achieve impressive high accuracy on coding benchmarks (Gunasekar et al., 2023; Li et al., 2023). These approaches typically require a powerful language model, such as GPT-3.5 and GPT-4 in (Eldan and Li, 2023), to synthesize new contents. Instead, (Maini et al., 2024) shows that compact models such as Qwen-1.8B and Mistral-7B are adequate to rephrase web data. This approach generates diverse, high-quality synthetic data that effectively lowers model perplexity and boosts performance across benchmarks. We adopt this main idea, but explore more prompts and show how to specialize them for low and high quality data.
@@ 5 Conclusion
For producing long-horizon pretraining tokens for LLMs from English Common Crawl data, we showed how to improve upon the state of the art and achieve better trade-offs between benchmark accuracy and data quantity, as measured by number of unique real tokens. Specifically, we showed the efficacy of ensembling model-based quality filters, rephrasing low and high quality documents, and reducing the reliance on non-learned heuristics. The dataset is public and split by quality level and type (actual data vs. different types of synthetic data), enabling the community to do further experiments on quality vs. diversity and how to build effective short and long horizon curricula.
@@ 6 Limitations
Some of the key limitations of our work are as follows. For the model-based filter ensembling and quality bucketing, we only had time and resources to try a single strategy. Though it is effective, it is possible this could be improved upon in future work, especially to improve the sensitivity at the higher-quality end of the spectrum. For the rephrased data, we did not verify the factual accuracy or fidelity to the original contents. More work is required to understand the risks of hallucinations or loss of content diversity in this setting and how to mitigate them. We also only looked at rephrasing low and high quality data. It could be interesting to explore how to best rephrase medium quality data as well. We did not do ablations on all parts of the pipeline. There is probably room for improvement with, for example, the language identification. Overall, we tried our methods only on English text. More work is needed to adapt our methods to other languages.
Finally, we did not decontaminate the dataset, as there is not yet a strong consensus on how to best do this and the impact is uncertain and debated, especially for large models trained over large token horizons. We note that the datasets we compare against (FineWeb-Edu, DCLM) were released without decontamination, and the model we compare against (Meta Llama 3.1) was also trained on contaminated data. DCLM reports some contamination analysis, but the findings suggest contamination is not a key factor: e.g., MMLU actually increases after decontamination, and DCLM does better than FineWeb on MMLU, even though FineWeb has more MMLU contamination (see Section 4.6 and Appendix N in Li et al. (2024)). Still, it would be interesting to better understand the impact of contamination for different model sizes and different token horizons, and we hope the community can explore such questions on this public dataset.